OpenStudy (anonymous):

Multiply the complex numbers (3+2i)(3-2i)

OpenStudy (anonymous):

i believe it is 13

OpenStudy (anonymous):

using foil

OpenStudy (anonymous):

9-6i+6i-4(-1) = 9+4 = 13

OpenStudy (anonymous):

i^2 = (-1)*

OpenStudy (anonymous):

and that is it worked out